Using 2013 rates, , the tax owed by each of the following people has been calculated assuming that they all claim the standard deduction and neglect any tax credits.

Deidre is single with no dependents. Her adjusted gross income is $80,000. Rounded to the nearest​ dollar, Deirdre's income tax is $13,429.

Robert is a head of household taking care of two dependent children. His adjusted gross income is $80,000. Rounded to the nearest​ dollar, his income tax is $9,340.

Jessica and Frank are married with no dependents. They file jointly. They each have $80,000 in adjusted gross income for a combined $160,000. Their combined income​ tax, rounded to the nearest​ dollar, is $26,666.

Complete parts​ (a)-(e) below.

a. Calculate the FICA tax owed by each of the three​ sets, assuming the given gross adjusted incomes came from ordinary wages. Note that the FICA tax is​ 7.65% on the first $113,700 of income from wages and​ 1.45% on any income from wages in excess of $113,700.

Deidre $___ Robert $___ Jessica and Frank $____   ​(Round to the nearest dollar as​ needed.)

b.

c

d

e
Answer #1

Diedre=gross wage*7.65%=80000*7.65%=6120

Robert=gross wage*7.65%=80000*7.65%=6120

Jessica & Frank =gross wage*7.54%=160000*7.65%=12240